# StratLearn-z: Improved photo-$z$ estimation from spectroscopic data subject to selection effects

### v1

@@ 4 StratLearn-z performance
To obtain more robust numerical results for the performance metrics we run 15 different realisations for each covariate shift scenario. The datasets for each realisation are obtained as follows:
  • the full dataset is randomly permuted;
  • rejection sampling as described in Sec. 3.1 is applied to the randomly permuted dataset, once per covariate shift scenario, resulting in four different source/target sets;
  • random sampling is applied to each set, so that source sets have 20,000 objects and target sets have 63,000 objects (62,000 for the weak CS scenario, see Sec. 3.1).
@@ 4.2 Conditional redshift distribution metrics
where $\hat{p}_{T}(z|x)$ is the conditional redshift distribution prediction for the target set given the observed photometry $x$. Under the the assumption that $\hat{p}_{T}(z|x)$ exactly represents the conditional distribution of the true redshift, $z_{\rm spec}$, given $x$, PIT is uniformly distributed over the unit interval. On the other hand, shifts from uniformity in the PIT distribution suggest systematic deviations in the estimates of the conditional distributions of redshift, while over-estimation (under-estimation) of the spread of the conditional distribution of redshift, i.e., the uncertainty associated with the estimated $\hat{z}_{\rm photo}$, appear as a bump (dip) in the histogram of PIT. The PIT distribution is commonly used to visually evaluate the quality of estimated photo-$z$ distributions (see e.g. Bordoloi et al., 2010; Polsterer et al., 2016; D’Isanto & Polsterer, 2018; Tanaka et al., 2018; Euclid Collaboration et al., 2020).
However, we note that the PIT distribution is not an indicator of the information content of the estimated conditional density. Indeed, Schmidt et al. (2020) shows that a method with no predictive power can produce a perfectly uniform PIT distribution. (See discussion of the TrainZ estimator, which assigns the marginal redshift distribution of the source data to each galaxy in the target set. This distribution does not depend on the photometric data and produces the same photo-$z$ estimate for every galaxy in the target set.)
@@ 4.3 Results and comparison to GPz
Figure 4: 2D histograms for the estimated $\hat{z}_{\rm photo}$ versus the true $z_{\rm spec}$ for the no CS case, $\alpha=1$, $\beta=1$. Left: StratLearn-z, right: GPz. The colorbar indicates the number of sources in each 2D bin.
Figure 5: Same as Fig. 4, but for the weak CS case, $\alpha=4$, $\beta=5$. Left: StratLearn-z, right: GPz.
Figure 6: Same as Fig. 4, but for the mild CS case, $\alpha=5$, $\beta=6$. Left: StratLearn-z, right: GPz.
Figure 7: Same as Fig. 4, but for the strong CS case, $\alpha=5$, $\beta=7$. Left: StratLearn-z, right: GPz.
@@
Table 2: Average performance metrics and the corresponding standard deviations for StratLearn-z and GPz for the four covariate shift scenarios considered.
@@
Figure 8: Violin plots showing the distribution of the metrics (averaged over 15 data realisations) for the four covariate shift scenarios. StratLearn-z results are shown in blue, while GPz results are shown in red. Black points with errorbars represent the mean and standard deviation for each case. There is a clear trend in the GPz results, which are strongly degraded by increasing covariate shift. On the other hand, the StratLearn-z results are only marginally impacted.
@@
We note that StratLearn-z outperforms GPz for all metrics and in all cases considered, even in the no CS case, which can likely be ascribed to better performance of the conditional density estimators. However, the difference is more pronounced with more covariate shift. Both the averaged RMSE and the normalised averaged RMSE, shown in the top panels of Fig. 8, show a similar trend: while for GPz they are nearly doubled when going from the case with no CS to the case with strong CS, they are only marginally impacted by the presence of covariate shift in the StratLearn-z results. Turning to the mean error, or bias (central left panel of Fig. 8), the impact of covariate shift on StratLearn-z leads to a mean error that is four times larger in the strong CS scenario with respect to the no CS case, while for GPz the increase is almost one order of magnitude. Moreover, we note that the averaged normalised error for StratLearn-z is always consistent with zero, except for the no CS case where it is slightly lower. Concerning the FR metric we find again that with StratLearn-z FR is consistent across covariate shift scenarios, with the FR15 (FR05) being only 0.5% (2.5%) worse in the strong CS case with respect to the no CS case. For GPz on the other hand FR15 decreases from $\sim 99$ in the no CS case to $\sim 96$ in the strong CS case. Similarly, the FR05 ranges from $\sim 98$ to $\sim 95$ for StratLearn-z, and from $\sim 95.5$ to $\sim 85$ for GPz.
@@
To evaluate the predicted conditional redshift distributions, Fig. 9 plots the histogram of PIT obtained from StratLearn-z and GPz for the four scenarios, focusing on one specific realisation. We checked that the PIT histograms are not significantly different across the 15 realisations. We see a significant accumulation of mass near 0.5 in the PIT distribution obtained from StratLearn-z, indicating an over-estimation of uncertainty in the conditional density estimates. However, the peak is centered at 0.5 and the distributions are symmetric, which suggests that most of the StratLearn-z conditional density estimates are well-centered around the true redshift values. Similar considerations hold for all covariate shift cases, highlighting once again the robustness of StratLearn-z results in presence of unrepresentative source datasets. On the other hand, the PIT distributions obtained from GPz are significantly flatter, but progressively less symmetric the stronger the covariate shift. We argue that the bump in the PIT distribution of StratLearn-z, although not ideal, is merely a symptom of our conditional redshift distribution estimates being too broad, i.e., too conservative (see also Appendix C for a toy model demonstration). Additionally, we notice that the peaks at 0 and 1 in the PIT distributions are substantially lower for StratLearn-z compared to GPz, signaling a smaller fraction of catastrophic outliers for StratLearn-z, and they are below 1, which is the value they should take if PITs were evaluated with the perfectly specified conditional distributions.
@@
Figure 9: PIT distribution for StratLearn-z (left) and GPz (right) for all cases, as detailed in the legend: in blue the no CS case, in green the weak CS case, in orange the mild CS case and in green the strong CS case.
@@ 5 Conclusions
To evaluate the quality of the predicted conditional redshift distribution we plot the histogram of the PIT, for each covariate shift scenario and for both codes. While StratLearn-z seems to be overestimating the spread of the conditional redshift distribution, the PIT distribution shows a reduced fraction of catastrophic outliers and symmetry, with a large bump around 0.5, i.e., the StratLearn-z conditional densities appear well-centered around the true redshift values. Despite the overestimation of photo-$z$ pdf spread, we note that the PIT distribution for StratLearn-z is also consistent across covariate shift scenarios, highlighting again the robustness of the method. One assumption made in the paper is that the covariate shift assumption holds, i.e., there are no unmeasured (confounding) covariates that are predictive for redshift, and also associated to the selection process into source/target set. Covariate shift is often assumed when estimating photo-$z$, e.g, (Izbicki et al., 2017; Freeman et al., 2017), but may not be completely accurate for real data, e.g., when quality cuts are applied to the spectroscopic sample (Hartley et al., 2020), which in turn do not cover the same color-magnitude space as the photometric observation. Investigation of such violations will be the topic of a future work.
@@
In summary, we found StratLearn-z provides stable predictions even in the most extreme scenario considered, in particular for high redshift regions where source/training data are lacking. Our analysis shows promising results, specifically concerning applications of the method to current and upcoming galaxy surveys, that will have to face the problem posed by incomplete and/or unrepresentative spectroscopic source sets to maximise the exploitation of photometric data.
@@
Acknowledgements: CM’s work is supported by the Fondazione ICSC, Spoke 3 Astrophysics and Cosmos Observations, National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) Project ID CN_00000013 “Italian Research Center on High-Performance Computing, Big Data and Quantum Computing” funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di "campioni nazionali di R&S (M4C2-19 )" - Next Generation EU (NGEU). RT acknowledges co-funding from Next Generation EU, in the context of the National Recovery and Resilience Plan, Investment PE1 – Project FAIR “Future Artificial Intelligence Research”. This resource was co-financed by the Next Generation EU [DM 1555 del 11.10.22]. RT is partially supported by the Fondazione ICSC, Spoke 3 “Astrophysics and Cosmos Observations”, Piano Nazionale di Ripresa e Resilienza Project ID CN00000013 “Italian Research Center on High-Performance Computing, Big Data and Quantum Computing” funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di “campioni nazionali di R&S (M4C2-19 )” - Next Generation EU (NGEU). David van Dyk and Maximilian Autenrieth acknowledge partial support from the UK Engineering and Physical Sciences Research Council [EP/W015080/1, EP/W522673/1] and from the Marie-Skodowska-Curie RISE [H2020-MSCA-RISE-2019-873089] Grant provided by the European Commission. RT and AM have been partially supported by the PRO3 Scuole Programme ‘DS4ASTRO’. This work used the Julia language and in particular the DataFrames.jl, CSV.jl, Statistics.jl, Distances, GLM.jl packages, and the Gnuplot.jl package for Julia to generate the plots.
@@ Appendix A Distributions for all photometric bands
Figure 10: Source and target normalised distributions of $u$-band magnitudes for the four covariate shift cases, as stated in each panel. Orange histograms represent the source dataset, while in blue we plot the target datasets.

### v2

@@ 4 StratLearn-z performance
To obtain more robust numerical results for the performance metrics, we run 15 independent realisations for each covariate shift scenario. For each realisation, the datasets are constructed as follows:
  • rejection sampling, as described in Sec. 3.1, is applied to the full dataset to generate four distinct source/target pairs corresponding to the different covariate shift scenarios;
  • from each of these sets, we randomly sample 20,000 objects for the source set and 63,000 for the target set (62,000 for the weak covariate shift scenario; see Sec. 3.1).
@@ 4.2 Conditional redshift distribution metrics
where $\hat{p}_{T}(z|x)$ is the conditional redshift distribution prediction for the target set given the observed photometry $x$. Under the assumption that $\hat{p}_{T}(z|x)$ exactly represents the conditional distribution of the true redshift, $z_{\rm spec}$, given $x$, PIT is uniformly distributed over the unit interval. On the other hand, shifts from uniformity in the PIT distribution suggest systematic deviations in the estimates of the conditional distributions of redshift, while over-estimation (under-estimation) of the spread of the conditional distribution of redshift, i.e., the uncertainty associated with the estimated $\hat{z}_{\rm photo}$, appear as a bump (dip) in the histogram of PIT. The PIT distribution is commonly used to visually evaluate the quality of estimated photo-$z$ distributions (see e.g. Bordoloi et al., 2010; Polsterer et al., 2016; D’Isanto & Polsterer, 2018; Tanaka et al., 2018; Euclid Collaboration et al., 2020).
However, we note that the PIT distribution is not an indicator of the information content of the estimated conditional density. Indeed, Schmidt et al. (2020) shows that a method with no predictive power can produce a perfectly uniform PIT distribution. As an example, the TrainZ estimator described in Schmidt et al. (2020) assigns the marginal redshift distribution of the source data to each galaxy in the target set. This distribution does not depend on the photometric data and produces the same photo-$z$ estimate for every galaxy in the target set, however, TrainZ is able to achieve a perfect PIT.
Additionally, we compute the continuous rank probability score (CRPS, Hersbach (2000); Polsterer et al. (2016)), defined as the integral of the squared difference between the redshift cumulative distribution function and the Heaviside step function centered at the true redshift value. More precisely, for each galaxy $i$, the CRPS is given by
${\rm CRPS}^{(i)}=\int_{-\infty}^{+\infty}\left[{\rm CDF}^{(i)}(z)-{\rm H}(z-z_{\rm spec}^{(i)})\right]^{2}\mathrm{d}z,$ (10)
where ${\rm CDF}^{(i)}(z)$ is the cumulative distribution function of the estimated conditional redshift distribution, and ${\rm H}(z-z_{\rm spec}^{(i)})$ is the Heaviside step function, i.e.,
${\rm H}(z-z_{\rm spec}^{(i)})=\begin{cases}0&\text{if }z<z_{\rm spec}^{(i)},\\ 1&\text{if }z\geq z_{\rm spec}^{(i)}.\end{cases}$ (11)
The CRPS quantifies the discrepancy between the estimated conditional redshift distribution and a Dirac delta function at $z_{\rm spec}^{(i)}$; it deviates from zero when the conditional distribution is broad or miscentered from the true redshift $z_{\rm spec}^{(i)}$. The median CRPS across all objects in the target dataset serves as a measure of the overall quality of the estimated redshift distributions.
@@ 4.3 Results and comparison to GPz
Figure 4: 2D histograms for the estimated $\hat{z}_{\rm photo}$ versus the true $z_{\rm spec}$ for the no covariate shift (CS) case, $\alpha=1$, $\beta=1$. Left: StratLearn-z, right: GPz. The colorbar indicates the number of sources in each 2D bin.
Figure 5: Same as Fig. 4, but for the weak covariate shift case, $\alpha=4$, $\beta=5$. Left: StratLearn-z, right: GPz.
Figure 6: Same as Fig. 4, but for the mild covariate shift case, $\alpha=5$, $\beta=6$. Left: StratLearn-z, right: GPz.
Figure 7: Same as Fig. 4, but for the strong covariate shift case, $\alpha=5$, $\beta=7$. Left: StratLearn-z, right: GPz.
@@
Table 2: Average performance metrics and the corresponding standard deviations for StratLearn-z and GPz for the four covariate shift (CS) scenarios considered.
@@
Figure 8: Violin plots showing the distribution of the metrics (averaged over 15 data realisations) for the four covariate shift (CS) scenarios. StratLearn-z results are shown in blue, while GPz results are shown in red. Black points with errorbars represent the mean and standard deviation for each case. There is a clear trend in the GPz results, which are strongly degraded by increasing covariate shift. On the other hand, the StratLearn-z results are only slightly impacted.
@@
We note that StratLearn-z outperforms GPz for all metrics and in all cases considered, even in the no covariate shift case, which can likely be ascribed to better performance of the conditional density estimators. However, the difference is more pronounced with more covariate shift. Both the averaged RMSE and the normalised averaged RMSE, shown in the top panels of Fig. 8, show a similar trend: while for GPz they are nearly doubled when going from the case with no covariate shift to the case with strong covariate shift, they are only marginally impacted by the presence of covariate shift in the StratLearn-z results. Turning to the mean error, or bias (central left panel of Fig. 8), the impact of covariate shift on StratLearn-z leads to a mean error that is four times larger in the strong covariate shift scenario with respect to the no covariate shift case, while for GPz the increase is almost one order of magnitude. Moreover, we note that the averaged normalised error for StratLearn-z is always consistent with zero, except for the no covariate shift case where it is slightly lower. Concerning the FR metric we find again that with StratLearn-z FR is consistent across covariate shift scenarios, with the FR15 (FR05) being only 0.5% (2.5%) worse in the strong covariate shift case with respect to the no covariate shift case. For GPz on the other hand FR15 decreases from $\sim 99$ in the no covariate shift case to $\sim 96$ in the strong covariate shift case. Similarly, the FR05 ranges from $\sim 98$ to $\sim 95$ for StratLearn-z, and from $\sim 95.5$ to $\sim 85$ for GPz.
@@
To evaluate the predicted conditional redshift distributions, Fig. 9 plots the histogram of PIT obtained from StratLearn-z and GPz for the four scenarios, focusing on one specific realisation. We checked that the PIT histograms are not significantly different across the 15 realisations. There is an accumulation of mass near 0.5 in the PIT distribution obtained from StratLearn-z, indicating an over-estimation of uncertainty in the conditional density estimates. This may prove to be an issue when high-precision estimates are needed, however, we note that the peak is centered at 0.5 and the distributions are symmetric, which suggests that most of the StratLearn-z conditional density estimates are well-centered around the true redshift values. Similar considerations hold for all covariate shift cases, highlighting once again the robustness of StratLearn-z results in presence of unrepresentative source datasets. On the other hand, the PIT distributions obtained from GPz are significantly flatter, but progressively less symmetric the stronger the covariate shift.
The mode in the PIT distribution of StratLearn-z indicated that our conditional redshift distribution estimates are too broad, i.e., too conservative (see also Appendix C for a toy model demonstration). To better understand this result, we inspected the PIT distribution obtained by replacing our conditional density estimators with GPz, i.e., by running GPz on the stratified data. The resulting PIT distributions are very similar to those obtained from GPz run on the unstratified data (i.e., the distribution shown in the right panel of Fig. 9), suggesting that the mode in the PIT obtained from StratLearn-z can be attributed to the conditional density estimators rather than the stratification. In future work, a possible improvement of the overall StratLearn-z pipeline could be to add GPz to our framework as an additional estimator to potentially flatten the PITs. We note that the covariate shirt scenarios considered here are relatively simple and performance as measured by PIT may vary with more realistic covariate shift scenarios.
@@
Figure 9: PIT distribution for StratLearn-z (left) and GPz (right) for all cases, as detailed in the legend: in blue the no covariate shift (CS) case, in green the weak covariate shift case, in orange the mild covariate shift case and in green the strong covariate shift case.
Table 3: Average of the median CRPS and its standard deviation over 15 realisations for the four covariate shift (CS) scenarios considered, comparing StratLearn-z and GPz (a smaller value indicates better performance).
StratLearn-z GPz
no CS $0.0079\pm 0.0001$ $0.0069\pm 0.0005$
weak CS $0.0101\pm 0.0006$ $0.0092\pm 0.0012$
mild CS $0.0105\pm 0.0007$ $0.0093\pm 0.0013$
strong CS $0.0112\pm 0.0010$ $0.0092\pm 0.0011$
Tab. 3 reports the median CRPS results for all four scenarios, both for StratLearn-z and GPz, averaged over all 15 realizations. GPz slightly outperforms StratLearn-z in all cases considered, which is to be expected given that GPz performs better than StratLearn-z as measured by PIT. However, we expect that the ability of StratLearn-z to better predict the mode of conditional redshift distributions (as reflected by its better performance in terms of the point estimate metrics reported in Table 2 and Fig. 8) should also be reflected in the CRPS, particularly at higher redshifts. To investigate the relative performance of the two methods in different redshift ranges, we computed the CRPS in ten redshift bins by taking the median CRPS in each bin (for each realisation), and then averaging over the realisations in each bin. The results are shown in Fig. 10 for the two extreme cases of no covariate shift (left) and strong covariate shift (right). Both cases show similar behavior: in low redshift bins GPz outperforms StratLearn-z due to its tighter variance in the predicted redshift distributions. However, at higher redshifts StratLearn-z consistently achieves a much lower CRPS, thanks to its improved prediction of the mode. The lower overall (i.e., across all redshifts) median CRPS of GPz is thus explained as the result of there being many more lower redshift objects (where GPz outperforms StratLearn-z) than higher redshift ones (where StratLearn-z vastly outperforms GPs). We also note that the CRPS obtained from StratLearn-z is roughly consistent across all redshifts, and shows a reduced dependence on the specific realisation than GPz, as demonstrated by StratLearn-z ’s smaller variance across realizations.
Figure 10: Binned CRPS as a function of redshift for no covariate shift (left) and strong covariate shift (right). Thin orange and light blue lines show the results for each realisation for GPz and StratLearn-z respectively, while thick lines with points represent the average over all realisations.
@@ 5 Conclusions
To evaluate the quality of the predicted conditional redshift distribution we plot the histogram of the PIT, for each covariate shift scenario and for both codes. While StratLearn-z overestimates the spread of the conditional redshift distribution, its PIT distribution shows a reduced fraction of catastrophic outliers and symmetry, with a large bump around 0.5, i.e., the StratLearn-z conditional densities appear well-centered around the true redshift values. Despite the overestimation of photo-$z$ pdf spread, we note that the PIT distribution for StratLearn-z is also consistent across covariate shift scenarios, highlighting again the robustness of the method. The GPz histograms are, however, more uniform than the StratLearn-z histograms, which suggests that incorporating GPz or other conditional density estimators within our StratLearn-z framework could lead to improved results. Additionally, we assessed the conditional redshift distributions via the CRPS metrics, which is informative on both on the variance of the redshift pdf and its mode with respect to the true redshift $z_{\rm spec}$. We find that, when comparing median CRPS computed over the whole redshift range, GPz slightly outperforms StratLearn-z for all covariate shift scenarios. However, when comparing the CRPS binned as a function of redshift, we showed that StratLearn-z consistently achieves lower values in the high redshift range, indicating once again that an optimal combination of conditional density estimators with our propensity score-based stratification procedure can further improve these results.
The fundamental assumption of this work is that the covariate shift condition holds true, i.e., that there are no unmeasured (confounding) covariates that are predictive for redshift, and also associated to the selection process into source/target set. Covariate shift is often assumed when estimating photo-$z$, e.g, (Izbicki et al., 2017; Freeman et al., 2017), but may not be completely accurate for real data. This is because the probability of obtaining a spectroscopic redshift may in fact depend on additional factors such as the type of object (e.g., passive galaxies are less likely to deliver a secure redshift compared to emission-line galaxies). Moreover, certain redshift ranges (such as the so-called ‘redshift desert’ $1.4\leq z\leq 2.5$) consistently lack reference redshifts, which occurs when the [O II] emission line shifts out of the optical range at approximately 0.9 µm, leaving a gap until strong UV lines redshift into the optical domain. Furthermore, spectroscopic redshifts measurements can in some cases deviate substantially from the true values, a situation known as ‘redshift failure’  (Hartley et al., 2020). To minimize the impact on subsequent analyses of poor quality redshifts, quality flags based on spectral features (e.g., S/N of emission and absorption lines, and the strength of the 4000 Å break) are used to identify and exclude unreliable redshifts (Hartley et al., 2020). Of course, such spectral features are only observed for spectroscopic confirmed galaxies and usually not available for galaxies in the photometric set, so they cannot be used for stratification. As discussed in Autenrieth et al. (2024a), the covariate shift condition remains valid if the spectroscopic covariates used for quality flagging do not provide additional information about redshift beyond what is already captured by the photometric covariates. Since the detailed procedure of obtaining secure redshift, as well as the definition of quality cuts are survey-specific, the definition and application of a general approach to this question is challenging. We will investigate potential violations of the covariate shift condition as well as potential correction strategies in a dedicated future study, focusing on topical and realistic survey selection procedures, e.g., as applied for Euclid (Laureijs et al., 2011; Euclid Collaboration et al., 2024).
@@
In summary, we found the propernsity-score based stratification procedure implemented in StratLearn-z provides stable predictions even in the most extreme scenario considered, in particular for high redshift regions where source/training data are lacking. Our analysis shows promising results, specifically concerning applications of the method to current and upcoming galaxy surveys, that will have to face the problem posed by incomplete and/or unrepresentative spectroscopic source sets to maximise the exploitation of photometric data.
@@
Acknowledgements: CM’s work is supported by the Fondazione ICSC, Spoke 3 Astrophysics and Cosmos Observations, National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) Project ID CN_00000013 “Italian Research Center on High-Performance Computing, Big Data and Quantum Computing” funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di "campioni nazionali di R&S (M4C2-19 )" - Next Generation EU (NGEU). RT acknowledges co-funding from Next Generation EU, in the context of the National Recovery and Resilience Plan, Investment PE1 – Project FAIR “Future Artificial Intelligence Research”. This resource was co-financed by the Next Generation EU [DM 1555 del 11.10.22]. RT is partially supported by the Fondazione ICSC, Spoke 3 “Astrophysics and Cosmos Observations”, Piano Nazionale di Ripresa e Resilienza Project ID CN00000013 “Italian Research Center on High-Performance Computing, Big Data and Quantum Computing” funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di “campioni nazionali di R&S (M4C2-19 )” - Next Generation EU (NGEU). DvD and MA acknowledge partial support from the UK Engineering and Physical Sciences Research Council [EP/W015080/1, EP/W522673/1] and from the Marie-Skodowska-Curie RISE [H2020-MSCA-RISE-2019-873089] Grant provided by the European Commission. MA also acknowledges partial support from the European Union’s Horizon 2020 research and innovation programme under European Research Council Grant Agreement No 101002652 (PI K. Mandel) and Marie Skłodowska-Curie Grant Agreement No 873089. RT and AM have been partially supported by the PRO3 Scuole Programme ‘DS4ASTRO’. This work used the Julia language and in particular the DataFrames.jl, CSV.jl, Statistics.jl, Distances, GLM.jl packages, and the Gnuplot.jl package for Julia to generate the plots.
@@ Appendix A Distributions for all photometric bands
Figure 11: Source and target normalised distributions of $u$-band magnitudes for the four covariate shift (CS) cases, as stated in each panel. Orange histograms represent the source dataset, while in blue we plot the target datasets.